# TuneNSearch: a hybrid transfer learning and local search approach for solving vehicle routing problems

### v2

@@ 2.3 Multi-task learning and transfer learning for VRPs
While these techniques have been widely studied in computer vision (Yuan et al., 2012) and natural language processing (Dong et al., 2019), their applications in combinatorial optimization remain relatively new. Recently, a few recent studies have begun exploring these methods in this domain, all utilizing reinforcement learning for training. Lin et al. (2024) proposed pre-training a backbone model on a standard TSP and subsequently fine-tuning to adapt to other routing variants, including the orienteering problem, the prize collecting TSP and CVRP. Their approach modified the neural network architecture of the pre-trained model by incorporating additional layers tailored to the unique constraints of each routing variant considered in the fine-tuning phase. Liu et al. (2024) modified the attention model to include an attribute composition block. This technique updates a problem-specific attribute vector, which dynamically activates or deactivates relevant problem features depending on the VRP variant being solved. The model includes four attributes that represent capacity constraints, open routes, time windows, and route limits. Essentially, it functions as a multi-task learning model trained on data from various VRP variants. Zhou et al. (2024a) aimed to improve generalization by incorporating a mixture-of-experts layer and a gating network. Specifically, the mixture-of-experts consists of multiple specialized sub-models, or “experts”, each one designed to handle different problem variants. The gating network then selects which experts to activate depending on the input, enabling the model to generalize to various tasks more effectively. Finally, Berto et al. (2025) introduced mixed batch training, a technique that enhances convergence by sampling multiple VRP variants within the same training batch. This allows the model to learn from multiple variants simultaneously at each optimization step, improving overall training efficiency.
@@
Despite these contributions, previous works have struggled to achieve strong performance. They are generally outperformed by specialized models trained for specific VRP variants, as well as other traditional optimization methods, such as OR-Tools. This raises a critical concern: if multi-task learning models exhibit a significant performance drop, their practicality in real-world applications becomes challenging to justify. Notably, the transfer learning approach proposed by Lin et al. (2024) achieved good performance. However, this came at the cost of extensive fine-tuning, requiring the same number of epochs as training a model from the beginning. Moreover, their method was limited to the orienteering problem, prize collecting TSP and CVRP, with no exploration of more prevalent VRP variants. Motivated by the aforementioned gaps, we introduce TuneNSearch, a novel transfer learning method pre-trained on MDVRP data and fine-tuned for different VRP variants. TuneNSearch differs from existing work in various aspects: First, to improve the encoding of VRP’s features, we integrate POMO with the residual E-GAT model. While the residual E-GAT model has shown improvements over the attention model, to the best of our knowledge, no prior work has combined it with POMO. We demonstrate that this combination enables a more effective encoding, making a better use of the multiple starting nodes introduced by POMO; Second, while we draw inspiration from Lin et al. (2024), we propose pre-training our model on the MDVRP — a significantly more complex problem than the TSP. This allows the model to learn richer features, facilitating a more effective knowledge transfer across different VRP variants. In addition to generalizing effectively to single-depot variants, TuneNSearch also demonstrates strong performance on various MDVRP variants, tasks that are neglected by current methods; Third, most existing neural-based approaches rely solely on machine learning, with little integration with optimization techniques (Mazyavkina et al., 2021). As a result, although neural-based methods are generally competitive for solving instances with up to 100 nodes, their generalization to larger instances remains limited. To address this, we incorporate an efficient local search algorithm after model inference, using a diverse set of search operators to iteratively refine solutions. This hybrid approach allows TuneNSearch to generalize more effectively than existing neural-based models across a range of VRP variants, in both randomly generated and benchmark instances. It achieves minimal integrality gaps while adding a small computational burden.
@@ 3 Preliminaries
In this section we first describe the formulation of the MDVRP and then introduce a brief overview of general neural-based methods for VRPs. After, we present other VRP variants featured in our work, along with their respective constraints.
@@ 3.1 MDVRP description
The MDVRP is an extension of the classical VRP, in which multiple depots are considered. This problem can be stated as follows: a set $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{m}\}$ of $m$ depots, and a set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$ of $n$ customers are given. Combined, these sets form a set of nodes $\mathcal{V}=\mathcal{C}\cup\mathcal{D}$, where the total number of nodes is $n+m=g$. The edge set =$\mathcal{E}=\left\{e_{ij}\right\},\ i,j\in\left\{0,\ldots,g\right\}$ represents the travel distances between all nodes in the problem. Each instance can be characterized by a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. A fleet of vehicles, each with a capacity $Q$, is dispatched from all depots to serve the customers. Each customer $c_{i}$ has a specific demand $\delta_{i}$ and must be visited exactly once by a single vehicle. Once a vehicle completes its route, it must return to its starting depot. The solution tour $\tau$ represents the sequence of nodes visited in the problem. It consists of multiple sub-tours, where each sub-tour corresponds to the set of nodes visited by an individual vehicle. In other words, $\tau$ captures the complete routing plan, breaking it down into distinct routes assigned to different vehicles. A solution is feasible as long as the capacity of each vehicle is not exceeded, and each customer is served exactly once. The objective is to find the optimal tour $\tau^{\ast}$ that minimizes the total distance traveled by all vehicles.
@@ 3.2 VRPs as a Markov decision process
Most existing neural-based models for VRPs use the reinforcement learning framework for training (Sutton and Barto, 1998). This approach can be formulated as a Markov decision process, which consists of the following key components:
State: The state is an observation received by the reinforcement learning agent which represents the current situation of the environment. At each timestep $t$, the state $s_{t}$ contains the embeddings of all node features, processed by an encoder, along with contextual information about the current partial solution tour.
Action: Upon receiving the state $s_{t}$, the agent selects the next action $a_{t}$ based on the current state. The action space contains all the nodes that can be added to the current partial solution. A masking mechanism is applied to ensure the feasibility of all solutions. This mechanism removes unfeasible nodes from the candidate actions space at each timestep.
@@
State transition: The state transition describes how the environment evolves as the agent picks actions. At each timestep t, the agent transitions from state $s_{t}$ to the next state $s_{t+1}$ based on the action $a_{t}$. The selected node is then added to the current partial solution, and other constraints are updated, including: the demand of the selected node is marked as fully served, the vehicle’s remaining capacity is reduced by the demand of the selected node, and the masking mechanism is updated to reflect these changes.
@@
Reward: The reward is a value used to evaluate the quality of solutions generated by the reinforcement learning agent, acting as a feedback signal to guide the learning process toward minimizing the objective function. At each timestep t, the agent receives a reward $r_{t}\ =\ -cost(s_{t+1},\ \ s_{t})$ which reflects the negative distance traveled between states $s_{t}$ and $s_{t+1}$. Once the entire tour is completed, the cumulative reward $R\ =\ -cost(\tau)$ represents the negative total distance traveled in the solution $\tau$. The agent’s objective is to maximize its total cumulative reward, which aligns with minimizing the total distance traveled.
@@
Policy: To maximize the total cumulative reward, the agent learns a policy, parametrized by an attention-based neural network (policy network) with parameters $\theta$. This policy is a function, or model, that maps states to actions. In essence, the agent learns a heuristic to determine how it should select the next action $a_{t}$ given the current state $s_{t}$, which is why neural-based methods are often referred to as neural heuristics. At each timestep $t$, the policy network takes as input the state $s_{t}$ and outputs the probabilities of visiting each node next. The agent then selects the node greedily (i.e., the node with the highest probability) or by sampling (choose an action stochastically based on the probabilities). This process continues until the full tour $\tau$ is constructed. The probability of constructing a tour $\tau$ can be expressed as $p_{\theta}(\tau|\mathcal{G})\ =\ \prod_{t=1}^{Z}p_{\theta}(a_{t}|\mathcal{G},\ a_{<t})$, where $a_{t}$ denotes the selected node, $a_{<t}$ the current partial solution and $Z$ is the maximum number of steps. To train the model, most works use the REINFORCE algorithm (Williams, 1992), which is explained in more detail in Section 4.2.
@@
Different from existing approaches, TuneNSearch initially pre-trains the policy network specifically on MDVRP data, which allows the model to establish a solid foundation of knowledge transferable to both single- and multi-depot variants. Then, for each VRP variant, the parameters from the pre-training phase are loaded into the model and a short fine-tuning phase is performed. In this phase, the model undergoes further training using data specific to each VRP variant. This fine-tuning phase is more efficient than training a new model from the beginning for each variant, as the model benefits from the knowledge acquired during pre-training.
@@ 3.3 VRP variants
[…]
@@ 4 Methodology
In this section, we formally describe our proposed method. Our work makes significant advances in two key areas: the development of a method with great generalization capability over different VRP variants, and the improvement of the inference process via the intersection of operations research with artificial intelligence methods. To achieve this, we do the following: i) Design of a novel model architecture combining POMO and the residual E-GAT encoder; ii) Implementation of a pre-training and fine-tuning framework designed to facilitate adaptation to the most common VRP variants; iii) Integration of a local search algorithm which applies different search operators to iteratively improve the solutions found by the neural-based model. The following sections detail each of these components.
@@ 4.1 Model architecture
Below we outline the architecture of TuneNSearch, which is built upon POMO. Our method incorporates the residual E-GAT in the encoder, which has previously shown performance improvements over the attention model (Lei et al., 2022). Unlike standard attention-based encoders that rely solely on node coordinates, the residual E-GAT extends the original graph attention network (Veličković et al., 2018) by incorporating the information of edges $e_{ij}\in\ \mathcal{E}\ \forall\ i,\ j\ \in\left\{0,\ \ldots,\ g\right\}$ and introducing residual connections between consecutive layers. These enhancements allow the model to better capture the information of graph structures, deriving efficient representations and more accurate attention coefficients. When paired with POMO’s multiple starting nodes sampling strategy, this enhanced encoding allows the model to evaluate diverse solution trajectories from different starting points with more precision. The distance-aware attention mechanism ensures that each rollout receives more contextually relevant information, helping the model converge to higher-quality solutions. To the best of our knowledge, this is the first time the residual E-GAT encoder is combined with POMO. Fig. 1 presents an illustration of the encoder-decoder structure of TuneNSearch. The model first encodes the features of depots, customers, and edge distances using a residual E-GAT. The resulting embeddings, along with contextual information about the partial solution tour, are then passed through a multi-head attention (MHA) layer. Finally, a single-head attention (SHA) layer, followed by a softmax function, calculates the probability of selecting each node next.
@@ 4.1.1 Encoder details
[…]
@@
After this initial transformation, both $x^{(0)}$ and $\hat{e}={\hat{e}_{11},\hat{e}_{12},...,\hat{e}_{gg}}$ are used as inputs to a residual E-GAT module with $L$ layers. Here, the attention coefficient $\alpha_{ij}^{l}$ indicates the influence of the features of the node indexed by $j$ on the node indexed by $i$, at layer $l\in\left\{1,2,...,L\right\}$. The coefficient $\alpha_{ij}^{l}$ can be calculated according to Equation 5:
@@
Then, a residual connection between consecutive layers is applied, in which the output at layer $l$, for node indexed by $i$, can be represented as indicated in Equation 6:
$\hat{x}_{i}^{(l)}=(\sum\limits_{k=0}^{g}\alpha_{ik}^{l}W_{2}^{l}x_{i}^{(l-1)})+x_{i}^{(l-1)}$ (6)
where $W_{2}^{l}$ is also a learnable matrix.
Lastly, after each E-GAT layer, the output also passes through two batch normalization and one fully connected layer. See Equations 7 and  8.
$\hat{x}_{i}=BatchNorm^{l}(x_{i}^{(l-1)}+\hat{x}_{i}^{(l)})$ (7)
$x_{i}^{(l)}=BatchNorm^{l}(\hat{x}_{i}+FC(\hat{x}_{i}))$ (8)

### v3

@@ 2.3 Multi-task learning and transfer learning for VRPs
While these techniques have been widely studied in computer vision (Yuan et al., 2012) and natural language processing (Dong et al., 2019), their applications in combinatorial optimization remain relatively new. Recently, a few recent studies have begun exploring these methods in this domain, all utilizing reinforcement learning for training. Lin et al. (2024) proposed pre-training a backbone model on a standard TSP and subsequently fine-tuning it to adapt to other routing variants, including the orienteering problem, the prize collecting TSP and CVRP. Their approach modified the neural network architecture of the pre-trained model by incorporating additional layers tailored to the unique constraints of each routing variant considered in the fine-tuning phase. Liu et al. (2024) modified the attention model to include an attribute composition block. This technique updates a problem-specific attribute vector, which dynamically activates or deactivates relevant problem features depending on the VRP variant being solved. The model includes four attributes that represent capacity constraints, open routes, time windows, and route limits. Essentially, it functions as a multi-task learning model trained on data from various VRP variants. Zhou et al. (2024a) aimed to improve generalization by incorporating a mixture-of-experts layer and a gating network. Specifically, the mixture-of-experts consists of multiple specialized sub-models, or “experts”, each one designed to handle different problem variants. The gating network then selects which experts to activate depending on the input, enabling the model to generalize to various tasks more effectively.
While these approaches improve flexibility and generalization, they still face some limitations. First, they often require substantial training resources (on the order of several days of compute time) and still fall short of the solution quality offered by specialized models or methods such as OR-Tools. Second, their generalization ability is often limited to single-depot VRP variants, which do not capture the operational complexity of more realistic logistics settings. The MDVRP, for example, is essential for modeling real-world logistics systems that involve multiple dispatch centers, such as urban distribution networks or decentralized supply chains, but remains underexplored in the context of neural combinatorial optimization.
@@
In response, we propose TuneNSearch, a new transfer learning method initially pre-trained on the MDVRP and fine-tuned for multiple VRP variants. Our approach distinguishes itself in three key ways: First, to improve the encoding of VRP’s features, we integrate POMO with the E-GAT encoder. While the residual E-GAT model has shown improvements over the attention model (Lei et al., 2022), to the best of our knowledge, no prior work has combined it with POMO. We demonstrate that this combination enables a more effective encoding, making a better use of the multiple starting nodes introduced by POMO; Second, while we draw inspiration from Lin et al. (2024), we propose pre-training our model on the MDVRP, a significantly more complex problem than the TSP. This allows the model to learn richer features, facilitating a more effective knowledge transfer across different VRP variants. On single-depot problems, TuneNSearch matches the performance of models trained on the CVRP, while performing substantially better on MDVRP variants; Third, most existing neural-based approaches rely solely on machine learning, with little integration with optimization techniques (Mazyavkina et al., 2021). As a result, although neural-based methods are generally competitive for solving instances with up to 100 nodes, their generalization to larger instances remains limited. To address this, we incorporate an efficient local search algorithm after model inference, using a diverse set of search operators to iteratively refine solutions. This hybrid approach allows TuneNSearch to generalize more effectively than existing neural-based models across a range of VRP variants, in both randomly generated and benchmark instances. It achieves much lower relative performance deviations compared to other neural-based methods, while incurring only a small computational overhead.
@@ 3 Preliminaries
In this section we first describe the formulation of the MDVRP, followed by a brief overview of how neural-based methods formulate VRPs as a Markov decision process. After, we present other VRP variants featured in our work, along with their respective constraints.
@@ 3.1 MDVRP description
The MDVRP is an extension of the classical VRP, in which multiple depots are considered. This problem can be stated as follows: a set $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{m}\}$ of $m$ depots, and a set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$ of $n$ customers are given. Combined, these sets form a set of nodes $\mathcal{V}=\mathcal{C}\cup\mathcal{D}$, where the total number of nodes is $n+m=g$. The edge set $\mathcal{E}=\left\{e_{ij}:i,j\in\mathcal{V},\ i\neq j,\ (i,j)\notin\mathcal{D}\times\mathcal{D}\right\}$ represents the travel distances between distinct nodes in the problem, excluding direct connections between depots. Each instance can be characterized by a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. A fleet of vehicles, each with a capacity $Q$, is dispatched from all depots to serve the customers. Each customer $c_{i}$ has a specific demand $\delta_{i}$ and must be visited exactly once by a single vehicle. Once a vehicle completes its route, it must return to its starting depot. The solution $\tau$ represents the sequence of nodes visited in the problem. It consists of multiple routes, where each route corresponds to the set of nodes visited by an individual vehicle. In other words, $\tau$ captures the complete routing plan, breaking it down into distinct routes assigned to different vehicles. A solution is feasible as long as the capacity of each vehicle is not exceeded, and each customer is served exactly once. The objective is to find the optimal solution $\tau^{\ast}$ that minimizes the total distance traveled by all vehicles.
@@ 3.2 VRPs as a Markov decision process
Most existing neural-based models for VRPs use the reinforcement learning framework for training (Sutton and Barto, 1998). In our approach, solutions are built one route at a time. At the start of each route, the agent selects a depot from which the vehicle will depart. The agent then starts to append customers to that route, and the route remains active until the agent decides to close it by selecting the starting depot again. In this way, the agent manages each route independently, ensuring that all assignments for one route are completed before initiating another one. Once a route is closed, a new route is initialized from a depot (either the same or a different one) chosen by the agent, and this process continues until all customers have been assigned. This approach can be formulated as a Markov decision process, which consists of the following key components:
State: The state is an observation received by the reinforcement learning agent which represents the current situation of the environment. Here, the environment is the system in which the agent operates, which in this case is a VRP instance. At each timestep $t$, the state $s_{t}$ includes embeddings of node features, which are vector representations capturing the relevant properties of each node in a continuous and numerical form, suitable to be processed by machine learning models. These embeddings are generated by an encoder, a neural network that transforms raw node features into an informative latent space. In addition to the node embeddings, $s_{t}$ also contains contextual information about the current partial solution, such as the vehicle’s remaining load, the total route length so far, and the embedding of the last visited node.
Action: Upon receiving the state $s_{t}$, the agent selects the next action $a_{t}$ based on the current state. Actions can be of two types:
  • 1.
    Depot node: chosen either to initialize a new route (by selecting the depot from which the vehicle departs) or to terminate the active route (by returning the vehicle to its starting depot).
  • 2.
    Customer node: chosen to extend the active route by visiting an unserved customer, subject to feasibility constraints.
To enforce feasibility, a masking mechanism is applied at each timestep to dynamically restrict the action space. For customers, this implies that nodes already visited and those whose demand exceed the remaining vehicle capacity are masked out from the action selection process. For depot nodes, availability depends on the stage of the solution. Before initializing a new route, all depots are available, because the agent must decide to which depot the next route will correspond to. At this point, customers cannot be selected. Once a depot is chosen, the feasible customers become available, along with the chosen depot itself (to allow the agent to finish a route whenever it deems appropriate). When a route is closed, all depots become available again, enabling the agent to decide where the next route will begin. We note that in non-MDVRP instances, where only a single depot exists, an action corresponding to a depot merely indicates the start or end of a route, as the depot itself is always fixed.
@@
State transition: The state transition describes how the environment evolves as the agent picks actions. At each timestep $t$, the agent transitions from state $s_{t}$ to the next state $s_{t+1}$ based on the action $a_{t}$. If a customer is selected, it is appended to the current active route. Its demand is marked as fully served, the vehicle’s remaining capacity is updated, and the mask excludes that customer from future actions. If the action corresponds to the depot that initiated the active route, the route is closed and the vehicle returns to its departing depot. If, however, the depot is selected at the start of a new route, it sets the route’s origin. In both cases, the masking mechanism is updated following the previously explained procedure.
@@
Reward: The reward is a scalar value that tells the agent how good or bad its action was in a given state, acting as a feedback signal to guide the learning process towards optimizing a specific objective. In the context of VRPs, at each timestep t, the agent receives a reward $r_{t}\ =\ -cost(s_{t+1},\ \ s_{t})$ which reflects the negative distance traveled between states $s_{t}$ and $s_{t+1}$. Once the entire solution is completed, the cumulative reward $R\ =\ -cost(\tau)$ represents the negative total distance traveled in $\tau$. The agent’s objective is to maximize its total cumulative reward, which aligns with minimizing the total distance traveled.
@@
Policy: To maximize the total cumulative reward, the agent learns a policy, parametrized by an attention-based neural network (policy network) with parameters $\theta$. This policy is a function, or model, that maps states to actions. In essence, the agent learns a heuristic to determine how it should select the next action $a_{t}$ given the current state $s_{t}$, which is why neural-based methods are often referred to as neural heuristics. At each timestep $t$, the policy network takes as input the state $s_{t}$ and outputs the probabilities of visiting each node next. The agent then selects the node greedily (i.e., the node with the highest probability) or by sampling (choose an action stochastically based on the probabilities). This process continues until the full solution $\tau$ is constructed. The probability of constructing a solution $\tau$ can be expressed as $p_{\theta}(\tau|\mathcal{G})\ =\ \prod_{t=1}^{Z}p_{\theta}(a_{t}|\mathcal{G},\ a_{<t})$, where $a_{t}$ denotes the selected node, $a_{<t}$ the current partial solution and $Z$ is the maximum number of steps. To train the model, most works use the REINFORCE algorithm (Williams, 1992), which is explained in more detail in Section 4.2.
@@
Different from existing approaches, TuneNSearch initially pre-trains the policy network specifically on MDVRP data, which allows the model to establish a solid foundation of knowledge transferable to both single- and multi-depot variants. Then, for each VRP variant, the parameters from the pre-training phase are loaded into the model and a short fine-tuning phase is performed. In this phase, the model undergoes further training using data specific to each VRP variant. This fine-tuning phase is more efficient than training a new model from the beginning for each variant, as the model benefits from previously acquired transferable knowledge.
@@ 3.3 VRP variants
The VRP variants solved by TuneNSearch are: i) CVRP: considers a single depot, in contrast to the multiple depots considered in MDVRP; ii) VRP with backhauls (VRPB): in the classic CVRP, a vehicle departs the depot loaded, and each time it visits a customer, it delivers goods, thereby decreasing its load as the route progresses. These customers are known as linehaul customers. In the VRPB, some customers require picking up goods (backhaul customers), which increase the vehicle’s load. In this paper, we consider a mixed VRPB, where linehaul and backhaul customers can be visited in any order, allowing a vehicle to alternate between deliveries and pickups within the same route. However, at all times, the vehicle’s load must remain within its maximum capacity limit. This condition implies, for example, that if a vehicle leaves the depot fully loaded to its maximum capacity, it cannot serve a backhaul customer as its first stop, since this would immediately exceed its maximum capacity $Q$; iii) VRP with duration limit (VRPL): in this variant, the length of each route cannot surpass a predefined threshold limit; iv) Open VRP (OVRP): in the OVRP, vehicles do not need to return to the depot after completing their route; v) VRP with time windows (VRPTW): in the VRPTW, each node has a designated time window, during which service must be made, as well as a service time, representing the time needed to complete service at that location; vi) Traveling salesman problem (TSP): the TSP is a simplified form of the CVRP that involves only a single route. In this problem, there is no depot, and nodes have no demands. Each node must be visited exactly once, and the vehicle (or salesman) must return to the starting node at the end of the route. We note that the constraints outlined above can also be applied to the MDVRP, instead of the CVRP. For more details on the generation of instances specific to all the described variants, we refer readers to B.
@@ 4 Methodology
In this section, we formally describe our proposed method, detailing each of the following components: i) Design of a novel model architecture combining POMO and the E-GAT encoder; ii) Implementation of a pre-training and fine-tuning framework designed to facilitate adaptation to the most common VRP variants; iii) Integration of a local search algorithm which applies different search operators to iteratively improve the solutions found by the neural-based model.
@@ 4.1 Model architecture
Below we outline the architecture of TuneNSearch, which is built upon POMO. Our method incorporates the E-GAT in the encoder, which has previously shown performance improvements over the attention model (Lei et al., 2022). Unlike standard attention-based encoders that rely solely on node coordinates, the E-GAT extends the original graph attention network (Veličković et al., 2018) by incorporating the information of edges $e_{ij}\in\ \mathcal{E}\ \forall\ i,\ j\ \in\left\{0,\ \ldots,\ g\right\}$. These enhancements allow the model to better capture the information of graph structures, deriving efficient representations and more accurate attention coefficients. When paired with POMO’s multiple starting nodes sampling strategy, this enhanced encoding allows the model to evaluate diverse solution trajectories from different starting points with more precision. The distance-aware attention mechanism ensures that each rollout receives more contextually relevant information, helping the model converge to higher-quality solutions. To the best of our knowledge, this is the first time the E-GAT encoder is combined with POMO. Fig. 1 presents an illustration of the encoder-decoder structure of TuneNSearch. The model first encodes the features of depots, customers, and edge distances using an E-GAT. The resulting embeddings, along with contextual information about the partial solution, are then passed through a multi-head attention (MHA) layer. Finally, a single-head attention (SHA) layer, followed by a softmax function, calculates the probability of selecting each node next.
@@ 4.1.1 Encoder details
Our encoder first embeds the features of all nodes in the problem to a $h_{x}$-dimensional vector space through a fully connected layer. Specifically, each depot $d_{j}\in\mathcal{D}$ is characterized by features $n_{j}^{d}$, which include its two-dimensional coordinates. On the other hand, each customer $c_{i}\in\mathcal{C}$ has features $n_{i}^{c}$, which not only contain its coordinates, but also the demand information and the early and late time windows. These features are passed through embedding layers separately for depots and customers. The result is a set of embedded vectors $\widehat{n}_{j}^{d}\ \in\ \mathbb{R}^{h_{x}}\ |j=1,...,m$} for depots and $\widehat{n}_{i}^{c}\ \in\ \mathbb{R}^{h_{x}}\ |i=1,...,n$} for customers, as shown in Equations 1 and 2. These vectors are then stacked to form $E^{d}\in\mathbb{R}^{m\times h_{x}}$ and $E^{c}\in\mathbb{R}^{n\times h_{x}}$. After, they are concatenated into $x^{\left(0\right)}\in\mathbb{R}^{g\ \times\ h_{x}}$, as demonstrated in Equation 3. Here, $x^{(0)}$ represents all nodes in the latent $h_{x}$-dimensional vector. We also embed the edge features, which represent the Euclidean travel distances $e_{ij},\ i,\ j\ \in\ \left\{1,\ 2,\ \dots,\ g\right\}$, into a $h_{e}$-dimensional vector space, as described in Equation 4.
@@
After this initial transformation, both $x^{(0)}$ and $\hat{e}={\hat{e}_{11},\hat{e}_{12},...,\hat{e}_{gg}}$ are used as inputs to an E-GAT module with $L$ layers. Here, the attention coefficient $\alpha_{ij}^{l}$ indicates the influence of the features of the node indexed by $j$ on the node indexed by $i$, at layer $l\in\left\{1,2,...,L\right\}$. The coefficient $\alpha_{ij}^{l}$ can be calculated according to Equation 5:
@@
Next, the node representations are updated using the weighted sum of neighboring nodes, followed by a residual connection and batch normalization:
$\tilde{x}_{i}^{(l)}=\text{BN}^{l}\left(x_{i}^{(l-1)}+\sum\limits_{k=0}^{g}\alpha_{ik}^{l}W_{2}^{l}x_{k}^{(l-1)}\right)$ (6)
where $W_{2}^{l}$ is another learnable matrix, and BN denotes batch normalization.
Finally, the output of this block is passed through a feed-forward layer, again followed by batch normalization:
$x_{i}^{(l)}=\text{BN}^{l}\left(\tilde{x}_{i}^{(l)}+\text{FF}^{l}(\tilde{x}_{i}^{(l)})\right)$ (7)
This process is repeated for all $L$ layers to obtain the final encoded node representations. In essence, the encoder generates node embeddings that capture the important features of each node, providing a compact summary of the entire problem instance. These embeddings are computed once by the encoder, before the decoder begins the solution construction process.